OpenStudy (anonymous):

What integer needs to be added to the expression to create a perfect square? x^2 + 30x A. 15 B. 30 C. 225 D. 900

Directrix (directrix):

Take half the coefficient of the linear x term (30) and square it. That is the integer that needs to be added.

OpenStudy (anonymous):

A

Directrix (directrix):

No. That is the half of the 30. Now, square 15 to get the integer you need to add. What is the square of 15?

Directrix (directrix):

15^2 = 15 * 15 = ?

OpenStudy (anonymous):

225

Directrix (directrix):

Yes. Let's check. x^2 + 30x + 225 = (x + 15)(x + 15) = (x + 15)^2 which is a perfect square trinomial.